shift summary for owl shift 07.20.99

Jim Graham (jagraham@hep.uchicago.edu)
Tue, 20 Jul 1999 07:00:20 -0500

Shift Summary for owl shift, 07.20.99

On Shift: JAG, JAWS

Hours of beam delivered at all intensities:7.0
Hours of beam delivered at nominal intensity:7.0
Hours of E832 data (special runs not included):6.5
Hours of special runs (laser scans, muon runs...):0

Accounting of lost time:Dart fresh start to clear level 3 processing
problems.

========================================================================
Runs ended this shift:
Run 13950 - 11.4M events - BEAM_832

Tapes brought to FCC this shift: QK*158-160, QKB054
QKCC07, QKN013, QKNL12, QKNL13, QKLA08, QKSM07

========================================================================
Accesses during shift: PSh finished replacing fuse in clock-fanout.

========================================================================
Overview of shift:
Generally smooth data-taking shift. A few minor problems for which
we decided not to wake the appropriate experts.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13950 1 00:09 Start 832 after access.
1 00:10 Load FPEDs at start of run due to changed cosmic
counter 109 (-30 counts).
15 00:30 disk /ktev4c goes full. This is the backup database
disk, and the databases are backed up here in a cron
job after midnight. Call Theo, and he had us delete
old backups.
54 01:30 NMSEED alarm goes off, Horizontal beam is off (-8.3
wires vs. -2.8 nominal). Julie calls MCR. MCR had
stopped giving beam to meson primary and didn't retune
after that.
59 Beam tuned to normal.
125 02:58 Missing FFFF word, stream 5, rwp 9 (crate 53?).
Kfrend.
173 04:04 Missing FFFF word, stream 5, rwp 9 (crate 53?).
Kfrend.
05:07 Do gas checklist.
05:08 Get SDAQ alarm. Seems the sw_sdaq.dbs database is
too big, and needs to get spooled off, and a new
one created. However, this one is not as big as the
one from July 8, and we are still writing to it, so
I assume we can go for a while, and we hold off on
paging the SDAQ experts.
05:15 Julie finds wire 6, pl 15 is inefficient at 0.92
+- 0.014 level. Since this is an edge wire, we
send mail to DC experts, but don't page.
224 05:26 Missing FFFF word, stream 5, rwp 9 (crate 53?).
Kfrend
05:40 Onsplit requests QKNL14. Mounted
05:44 Run ends on full tapes.
05:50 Onsplit requests QKPE07. Mounted

13591 1 05:55 Start run.
06:01 End run because L3 processes look messed up; ktev1,2
CPU low, ktev4 high

13592 1 06:13 Drt_fresh_start. Overwrite tapes. L3 processes look
good now. In the process, pick up a new version of
the daq monitoring tool (Jim's hack). I was trying
to get the "% used" to show drive 4, since this drive
fills up the most quickly. However, I modified too
much, and now all drives show drive 4's usage. I've
fixed this, but it won't get picked up until the
next fresh_start.

06:30 Replace Blue Ice on alcohol monitor.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift
o Keep an eye on RF-button in SDAQ monitoring. If it goes red, call
Shanahan.
o Contact Theo/Ashkan about sw_sdaq.dbs database.
o Check DC wire 6, plane 15 for inefficiency.

Regards, Jim, for Julie